OpenStudy (anonymous):

HELP HELP PLEASE! Three points are on a coordinate plane: A(1, 5), B(-2, -4), and C(6, -4). 1. Write an equation in point-slope form of the line with slope -1 that contains point C. 2. Write an equation in point-slope form of the line that contains points A an B. 3. Write an equation of the line that contains B and C.

OpenStudy (saifoo.khan):

i can help!

OpenStudy (anonymous):

YAY! <3

OpenStudy (anonymous):

1. y = -x + c substitute (6,-4): -4 = -6 + c so c = 2 so y = -x + 2 2. slope = (-4 - 5)/(-2 - 1) = -9/-3 = 3 so y = 3x + c substitute (1,5): 5 = 3 + c so c = 2 so y = 3x + 2 3. slope = (-4 - -4)/(6 - -2) = 0 hence y = -4

OpenStudy (saifoo.khan):

1 = find the y-intercept 1st

OpenStudy (saifoo.khan):

lol, he solvd.

OpenStudy (anonymous):

lol ill still give u both medal s thx alot <3